# Towards Aligning Multimodal LLMs with Human Experts: A Focus on Parent–Child Interaction

### v1

@@ 5.2. Partial Alignment at the Level of Judgement
Overall alignment scores at the judgement stage were modest, reflecting the difficulty of capturing interpretive nuances. Nevertheless, we found that many-shot prompting consistently outperformed zero-shot, suggesting that incorporating examples grounded in SLP expertise helps models approximate expert reasoning more closely. This indicates that expert knowledge does play a role in supporting judgement-level alignment, even if the gains are incremental.
The results highlight the need for more deliberate strategies to improve judgement-level alignment. Few-shot prompting offers one pathway, but more sustained approaches such as fine-tuning with expert-labelled data or retrieval-augmented generation (RAG) may be necessary to embed richer interpretive cues. At the same time, because judgement involves a degree of subjectivity, future systems may need to actively explore how different forms of expertise shape interpretation, and provide mechanisms to surface or adapt to these perspectives rather than converge on a single fixed standard.
This raises a broader question about how to handle inherently subjective layers of expertise in human-AI alignment. At the observational level, alignment is more attainable because experts tend to converge on concrete cues such as gaze, action, or vocalisation. At the judgement level, however, interpretations vary and are often context-dependent, making consensus elusive. This challenge is not unique to joint attention: expert decision-making in many fields is notoriously tacit and difficult to distil into explicit rules, drawing on years of situated practice and subtle interpretive skill. For instance, in medicine, radiologists may differ in judging whether a faint shadow on a scan indicates pathology or normal variation. In law, judges interpreting similar facts may reach different conclusions based on contextual reasoning. In Parent-Child Interaction analysis, where interpretive diversity is intrinsic, the goal may be less about maximising consensus and more about enabling systems to represent, contrast, and adapt to multiple plausible viewpoints. This perspective opens up space for HCI and AI research to treat subjectivity not as noise to be eliminated, but as a critical design consideration in the development of judgment-sensitive systems.
In education, a raised hand may be judged by one teacher as evidence of engagement, while another views it as disruptive if out of turn. In UX evaluation, a pause may be seen by one expert as a usability breakdown, but by another as part of a normal learning curve. In creative work critiques, the same bold use of colour may be praised by one reviewer as innovative, yet criticised by another as unbalanced. These examples highlight that while observation alignment is attainable, judgement alignment may be less realistic or even undesirable. For exploratory domains like Parent-Child Interaction analysis, acknowledging and surfacing disagreement may be more valuable than enforcing consensus, as it allows users to reflect on alternative interpretations rather than rely on a single definitive output.
@@ 5.3. Limitations
The dataset of 25 YouTube videos was skewed toward short, likely neurotypical interactions, introducing bias and constraining generalisability. Our contribution is exploratory, and we used demonstration-based recordings as a pragmatic choice given the difficulty of collecting authentic, neurodiverse, and naturally occurring Parent-Child Interaction data. However, this bias led to very few segments labelled as Poor, meaning our analysis primarily reflects alignment in the Strong category and provides only partial evidence of model–expert agreement overall. Notably, one SLP emphasised that detecting Poor joint attention is often clinically more critical, and suggested that recordings featuring neurodiverse children (e.g., autism-focused datasets) may not only be more representative of practice but also easier to annotate consistently. Expanding future datasets to include longer interactions, neurodiverse populations (benton2014diversity), and recordings from real-world home or school settings will be essential for improving ecological validity.
@@ 6. Conclusion
This paper presented an exploratory study of alignment in the context of parent–child interaction analysis. By comparing three SLP’ perspectives on joint attention with MLLM outputs, we identified a clear contrast: observation-level alignment was relatively robust, while judgement-level alignment remained elusive due to differences among experts themselves. These findings suggest that MLLMs can serve as reliable observers when scaffolded by shared behavioural cues, but that interpretive alignment requires more nuanced approaches and may benefit from exposing diverse perspectives rather than enforcing consensus.

### v2

@@ 5.2. Partial Alignment at the Level of Judgement
Alignment at the judgement stage was modest, reflecting the difficulty of capturing interpretive nuances. Many-shot prompting consistently outperformed zero-shot prompting, suggesting that examples grounded in SLP expertise help the model approximate expert judging more closely. These gains were incremental, however, indicating that judgement-level alignment is shaped not only by prompting strategies but also by natural variation in expert perspectives.
Judgement in Parent-Child Interaction analysis is inherently subjective and context dependent. Different SLPs weight behavioural cues differently, and their interpretations can diverge even when they observe the same interaction. Similar variation also appears in other expert domains. Teachers may interpret a raised hand as engagement, whereas others see it as disruption; UX evaluators may view a pause as confusion, while others treat it as normal exploration; and reviewers in creative critique often make contrasting assessments of the same work. These examples illustrate that judgement alignment is fundamentally different from observation alignment.
In our dataset, heterogeneous annotations made it difficult to derive a single, coherent judgement space. We also observed pockets of higher agreement among certain experts, such as SLP1 and SLP2 ($\alpha=0.31$). When data are limited and expert perspectives vary widely, such internally consistent subgroups point to a possible direction for future work: examining whether alignment strategies that draw on individual experts or small groups could offer more stable early-stage models.
Design Takeaway #2: Calibrate AI Judgement to Task Needs Through Expert-Aligned Adaptation Given the heterogeneous annotation landscape, a practical direction for HCI systems is to support personalised, expert-aligned adaptation. For SLP-facing tools, tailoring the AI to an individual clinician’s own judgement tendencies may be particularly useful, as their criteria for a given task tend to remain relatively stable over time. For parent-facing tools, however, a consensus-informed representation derived from multiple experts may provide clearer and more balanced guidance. Developing such a system responsibly would require substantially larger and more diverse datasets, including naturalistic parent–child interactions and broader SLP annotations, to ensure that any consensus model is reliable, generalisable, and appropriate for use in parent-oriented contexts.
At the presentation level, SLPs also highlighted the importance of conveying joint attention as an overall pattern rather than a sequence of isolated segments. SLP3 suggested a colour-bar overview where each segment’s label is encoded as a colour across the entire timeline. Such visual summaries provide both SLPs and parents with an immediate, intuitive understanding of interaction quality and help them focus efficiently on segments that may require further attention.
Across these use cases, AI outputs should be treated as supportive references rather than definitive judgements. From a reliability standpoint, the final assessment should still be made by an SLP. Judgement in Parent-Child Interaction is only one part of an SLP’s broader role; clinicians also provide therapy, coach parents, interact with children, and tailor guidance to highly specific family contexts. These relational and instructional responsibilities cannot be automated. AI systems can, however, enhance clinical efficiency by surfacing behavioural patterns, organising information, and providing structured summaries that clinicians can interpret and incorporate into therapeutic practice. Future work should investigate how personalised expert aligned models and consensus informed models might be integrated into broader clinical workflows and how these two modes could support different stakeholders within real developmental assessment and intervention pathways.
@@ 5.3. Limitations
The dataset of 25 YouTube videos was skewed toward short, likely neurotypical interactions, introducing bias and constraining generalisability. Our contribution is exploratory, and we used demonstration-based recordings as a pragmatic choice given the difficulty of collecting authentic, neurodiverse, and naturally occurring Parent-Child Interaction data. However, this bias led to very few segments labelled as Poor, meaning our analysis primarily reflects alignment in the Strong category and provides only partial evidence of model–expert agreement overall. Notably, one SLP emphasised that detecting Poor joint attention is often clinically more critical, and suggested that recordings featuring neurodiverse children (e.g., autism-focused datasets) may not only be more representative of practice but also easier to annotate consistently. Expanding future datasets to include longer interactions, neurodiverse populations (Benton et al., 2014), and recordings from real-world home or school settings will be essential for improving ecological validity.
Human Editing
A methodological limitation of our approach is that the input to the Stage 2 judge process was based on manually corrected descriptions rather than the raw outputs produced by the MLLM in Stage 1. Although fewer than 15% of all entries required correction and each modification followed a minimal-editing principle that preserved the subject–verb–object structure, these corrections removed factual errors that would otherwise propagate into Stage 2. Our intention was to isolate the effect of the judge process itself, but this design choice means that the Stage 2 results likely represent an upper bound on performance in a fully end-to-end setting. Future work will evaluate the judge process using uncorrected model outputs to obtain a more conservative estimate of system behaviour.
@@ 6. Conclusion
This paper presented an exploratory study of alignment in the context of parent–child interaction analysis. By comparing three SLP’ perspectives on joint attention with MLLM outputs, we identified a clear contrast: observation-level alignment was relatively robust, while judgement-level alignment remained elusive due to differences among experts themselves. These findings suggest that MLLMs can serve as reliable observers when scaffolded by shared behavioural cues, and they highlight how future HCI systems may draw value from both individual or subgroup SLP judgements and areas of broader consensus across the expert cohort.